OpenStudy (anonymous):

Find the equation of a line given the point and slope below. Arrange your answer in the form y = mx + b, where b is the constant. 0,0) m=2

OpenStudy (anonymous):

equation of a straight line passing through (a,b) with slope m is given by \[y-a=m(x-b)\] upon simplifying you get y=mx-(mb+a)